What is the kinetic energy of a 1,500-kg car traveling at 110 kph if the car can reach this velocity in 12 seconds starting from rest? What is the power output of the engine in HP?

1
Expert's answer
2022-06-09T09:59:40-0400

Given:

"m=1500\:\rm kg"

"v=110\:\rm km/h=30.6\:\rm m/s"

"t=12\:\rm s"


"P=\frac{W}{t}=\frac{\Delta E}{t}=\frac{mv^2}{2t}"

"P=\frac{1500*30.6^2}{2*12}=5.84*10^4\:\rm W=78.3\:hp"